Based on the following partial financial statement, calculate the fixed asset newness ratio: Provide your answer in decimal form, with three significant digits.

Current Assets
Cash and Securities $2,069,775
Accounts Receivables $4,466,680.7
Retainage $955,878.57
Inventory $87,437.73
Earnings in Excess of Billings $286,350.52
Other Current Assets $208,891.83
Fixed Assets
Property $23,000,000
Construction Equipment $3,571,453.8
Vehicles $562,741.32
Office Equipment $67,102
Depreciation $-3,254,652.41
Other Fixed Assets $245,049.8
Answer #1

Fixed assets newness ratio = Net Fixed assets/total fixed assets (excluding land)

Gross Fixed assets=(23000000+3571453.8+562741.32+67102+245049.8) =

27446346.92

Net Fixed assets= Gross fixed assets-Depreciation

Fixed assets newness ratio= (27446346.92-3254652.41)/ 27446346.92

= 0.881